OpenStudy (anonymous):

24,673.32=12,000(1+(.08/4))^(4*t) can someone help me solve for t (time)?

OpenStudy (anonymous):

@NotTim @ghass1978 @campbell_st @ash2326 @AccessDenied

OpenStudy (campbell_st):

if the equation is \[24 673.32 = 12000(1 + (\frac{0.8}{4}))^{4t}\] ok... so divide both sides of the equation by 12000 or \[2.05611 = (1 + (0.08/4))^{4t}\] next take the log of both sides. Base e is fine\[\ln(2.05611) = 4t \times \ln(1 + (\frac{0.08}{4})\] this should make it easier to find t

OpenStudy (campbell_st):

you will need to know a few log laws to simplify it

OpenStudy (anonymous):

anyone in 11th at c academy

hero (hero):

^ no

hero (hero):

@mariomintchev, I already explained to you how to do it.

OpenStudy (anonymous):

ok im getting a weird answer

OpenStudy (anonymous):

i got like -4.6

OpenStudy (campbell_st):

well I got a sensible answer... its positive... which is important as I think t stands for time...

hero (hero):

You already know -4.6 definitely isn't it

OpenStudy (campbell_st):

try doing some simplification... 1 + 0.08/4 = 1.02 \[\frac{\ln(2.05611)}{\ln(1.02)} = 4t\]

OpenStudy (anonymous):

9.1

OpenStudy (campbell_st):

well its positive ... like time... so go and substitute it into the original equation and see if you get 24673.32.... best way of checking

OpenStudy (anonymous):

its correct :)

OpenStudy (campbell_st):

well then its a good solution.